# Examining the Source of Defects from a Mechanical Perspective for 3D Anomaly Detection

### v1

@@ 4.2 Corrective Force Prediction Network
Based on the motivation in Section 2, we propose the Corrective Force Prediction Network (CFP-Net), which estimates internal and external corrective forces. Since defective forces constitute continuous vector fields in physical space, our differentiable U-Net architecture explicitly preserves this continuity through its topology. By U-Net’s universal approximation capability for inverse physical processes, this design directly enables learning implicit defective forces from point clouds and predicting correction forces through rigorous inverse modeling, where its differentiable operators inherently maintain continuous force field embeddings. More analysis of CTP-Net to generate damaging forces and corrective forces is reported in Appendix B.3.
@@
where three convolutional layers project $C_{1}$-dimensional features into isomorphic spaces with preserved dimensionality but distinct representations, aligning with the target feature domain.
@@
$F_{C}=MLPs(F_{9}^{\prime}).\ F_{C}\in\mathbb{R}^{n\times 6}.$ (9)
@@
Anomaly Score during Inference. The norm of point-level corrective resultant force $F_{C}$ is directly considered as an anomaly score during the inference phase. This is attributed to Definition A2 in Section 2 that the model gives each point a corrective force that is a zero vector when the point is normal and a non-zero vector when it is abnormal, as described in Eq. 4. The maximum value of the point-level score is considered to be the sample-level score.
4.3 Prediction Loss
To enhance the reconstruction accuracy of corrective force fields, we propose a composite loss function that integrates three components: a symmetry loss derived from the physical constraint in Eq. 4 requiring opposing directions between internal and external correction forces in anomalous regions, a distance loss and a directional consistency loss to jointly optimize global restoration performance. The combinatorial loss function $\mathcal{L}_{comb}$ is defined:
$\mathcal{L}_{comb}$ $=\mathcal{L}_{dist}+\mathcal{L}_{dir}+\mathcal{L}_{sym},$ (10)
$=\frac{1}{N}\sum_{i=1}^{N}\bigg{(}\frac{\|F^{\prime}_{I,i}-F^{\prime}_{E,i}\|_{1}}{\|F^{\prime}_{I,i}\|_{2}+\|F^{\prime}_{E,i}\|_{2}+\epsilon}-\frac{F^{\prime}_{I,i}}{\|F^{\prime}_{I,i}\|_{2}+\epsilon}\cdot\frac{F^{\prime}_{E,i}}{\|F^{\prime}_{E,i}\|_{2}+\epsilon}\bigg{)}$
$+\frac{1}{N}\sum_{i=1}^{N}\left\|F_{D,i}-F_{C,i}\right\|-\frac{1}{N}\sum_{i=1}^{N}\frac{F_{D,i}}{\left\|F_{D,i}\right\|_{2}+\epsilon}\cdot\frac{F_{C,i}}{\left\|F_{C,i}\right\|_{2}+\epsilon}.$
@@ 4.3 Prediction Loss
Quality Control (QC) under cost limitations is an important topic in real industrial scenarios qc . Enterprises often encounter the following two scenarios in real manufacturing: S1: Multi-level QC is an essential part of the QC landscape, and it is a common strategy to use coarse QC to detect significantly normal samples and then fine QC to identify anomalies in subtle samples. S2: Moreover, the number of different subcategories of products produced for the same series of products can be changed by the economics of market demand. Using an assembly line to produce different subcategories of products is a common strategy within cost limitations.
To further simulate the real industrial manufacturing, we first propose a hierarchical quality control (HQC) based on a three-way decision by simplifying our model to form a pruned version, which is collocated with our original method to realize the multilevel quality control mentioned in S1. In addition, a dataset named Anomaly-IntraVariance with multiple subcategories is proposed for evaluating the performance of the model in the S2. Information on the dataset is briefly described in Section 5.1 and discussed in detail in the Appendix D.
@@
$\begin{cases}\mathcal{D}_{\text{normal}}=\{P_{i}\mid\text{rank}(s_{i})\leq bN,\ s_{i}=V_{P}(P_{i})\},\\ \mathcal{D}_{\text{test}}=V_{O}(\{P_{i}\}\setminus\mathcal{D}_{\text{normal}}).\end{cases}$ (11)
@@ 5.1 Implementation
Datasets. (1) The Anomaly-ShapeNet Li_2024_CVPR has over 1,600 positive and negative samples from 40 categories, leading to a more challenging setting due to the large number of categories. (2) The Anomaly-ShapeNet-New Li_2024_CVPR provides over 450 positive and negative samples crossing 12 categories, which is not been used as an evaluation dataset before. (3) The MvTec3D-AD Bergmann_2022 includes 4,147 RGB-D sample pairs from 10 categories, of which 894 are anomalous. We use the depth sample for detection evaluation. (4) The Real3D-AD liu2023real3d dataset consists of 1,254 large-scale, high-resolution point cloud samples from 12 categories, with the training set for each category only containing 4 normal samples. (5) We introduce the Anomaly-IntraVariance dataset, the first 3D dataset for simulating real-world production situations where an assembly line produces multiple products, and contains more than 1,000 samples from 16 categories. We report more details in the Appendix D.
Baselines. We selected BTF BTF , PatchCore 9879738 , M3DM m3dm , CPMF cao2023CPMF , Reg3D-AD liu2023real3d , IMRNet Li_2024_CVPR , R3D-AD r3dad , ISMP ismp and PO3AD PO3AD for comparison. We evaluate the excellent performance of these models on Anomaly-ShapeNet-New. We report experimental Details, how we chose baselines, and how we implemented baselines in Appendix A.1, A.2, and A.3, respectively.
@@
Table 1: O-AUROC performance of different methods on Anomaly-ShapeNet, where best and second-place results are highlighted in red and blue, respectively. The full version is in Appendix G
O-AUROC
@@
5.2 Comparison Experiment
Comparison on Anomaly-ShapeNet. Tables 10 and 11 provide the detection and segmentation results of different methods on Anomaly-ShapeNet across 40 categories, respectively. The proposed approach obtained an outstanding average O-AUROC and P-AUROC of 90.9% and 91.0%, outperforming the second-best methods by 7.0% and 1.2%, respectively. Moreover, our method reaches 94.0% of the average results and outperforms the second-best methods by 5.9%. Notably, our method achieved the best average ranking on O-AUROC, P-AUROC, and O-AURP, demonstrating that our approach outperforms our competitors in most categories.
Comparison on Anomaly-ShapeNet-New. Tables 14 and 13 provide the detection and segmentation results of different methods on Anomaly-ShapeNet-New across 12 categories. The MC4AD obtained an outstanding average O-AUROC and P-AUROC of 88.8% and 93.7%, respectively.
@@ 5.2 Comparison Experiment
Comparison on Real3D-AD. Tables 18 and 17 present the detection and segmentation results of different methods on Real3D-AD across 12 categories, respectively. Our model obtained O-AUROC and P-AUROC of 78.6% and 83.7%, respectively, maintaining excellent segmentation performance while achieving the detection performance of SOTA. This may be attributed that the model predicts better when the training set and testing set vary significantly.
@@
O-AUROC 0.652 0.678 0.702 0.764 0.754 0.786
P-AUROC 0.688 0.705 0.733 0.812 0.792 0.837
Method O-AUROC P-AUROC FPS
$w/o\ GMR$ 0.733 0.740 17.1
@@
Comparison on Anomaly-IntraVariance. Tables 19 present the detection and segmentation results of different methods on Anomaly-IntraVariance across 16 categories, where Group 1 includes two subspecies and Group 2 has four subspecies. Our model achieves 76.1% and 80.7% on Group 1 and obtains 62.5% and 67.8% on the more challenging Group 2. The average abnormal detection metrics performed 4.40% and 2.65% better than the second-best method.
5.3 Ablation and Parameter Sensitivity Experiments
@@
Analysis of multiple loss functions effectiveness.$V_{1}$” is to predict a point cloud that lacks symmetry constraints, resulting in a predicted imbalance of internal and external Corrective forces. Without $\mathcal{L}_{sym}$, the O-AUROC and P-AUROC decreased by 13.4% and 14.9%, respectively, which may be attributed to the force imbalance resulting in an incorrect offset to the normal point. Moreover, we designed “$V_{2}$”, where the model is supervised without $\mathcal{L}_{dir}$. The performance causes a degradation of the 10.8% and 13.2% anomaly detection metrics because the direction of the incorrect force is not constrained, resulting in the direction of the offset not being taken into account in the prediction of the offset. Analogous to “$V_{2}$”, “$V_{3}$” is proposed for constraining the correct size of the offset. The model performs 8.4% and 10.4% decrease in O-AUROC and P-AUROC, respectively, which may be attributed that the model meets challenges to learn the offset distance for both normal and pseudo-abnormal points without $\mathcal{L}_{dist}$. Therefore, all the proposed loss functions are crucial for predicting the incorrect force and thereby correctly predicting the corrective forces.
Figure 3: Qualitative results of localization.
@@
Analysis of modules effectiveness. In “$V_{4}$” it is observed that O-AUROC and P-AUROC are reduced by 2.2% and 2.5%, respectively, due to the replacement of our proposed anomaly generation with Norm-AS in PO3AD, which decreases the learned diversity of anomalies, and thus leads to a decrease in anomaly detection performance. In addition, our pruned version of “$V_{P}$” acquires a more efficient 20 FPS (>14.8 FPS) while maintaining excellent anomaly detection performance, making our DMD strategy potentially possible, as discussed in Section 5.4. Moreover, the model without GMR ($w/o\ GMR$) modules failed to capture the complementary features of external and internal Corrective forces and was not able to generate the resultant Corrective force, resulting in a decrease in the anomaly detection metrics by 5.3% and 9.7%, respectively. In contrast, the model directly outputs 3-channel offsets ($Out=3$) equivalent to the direct learning of the Corrective force, resulting in a decrease of 4.0% and 4.2% in P-AUROC and O-AUROC, respectively. More discussion about the use of the Corrective force is reported in Section 5.5. We finally evaluated the effectiveness of each layer and observed that the number of layers was reduced from 8 to 2, and 5.8% and 12.4% reduced the positive correlation of anomaly detection metrics. To balance performance and efficiency, we choose 8 layers for the experiments. More parameter sensitivity experiments are reported in the Appendix E.
Table 3: Comparison between One Model and our DMD.
@@
One Model DMD One Model DMD
O-AUROC P-AUROC FPS O-AUROC P-AUROC FPS O-AUROC P-AUROC FPS O-AUROC P-AUROC FPS
@@
5.4 Generalizability Experiments and Resource Analysis
Generalizability Experiments. We apply MC-Skip to the time-tested classical model, including PointNet pointnet , PointNet++ pointnet2 , PointTransformer pointtransformer , and DGCNN dgcnn , to validate the generalizability of our model. As shown in Table 4(a), the accuracies with and without normal participation on the classification task wu20153dshapenetsdeeprepresentation increased on average by 0.65% and 0.6%, respectively, and the instance average IoU and class average IoU on the segmentation task shapenet increased on average by 0.775% and 0.675%, respectively. Moreover, our DMD strategy performs well on existing 3D anomaly detection models and generalizes to memory bank, as reported in Table 3. Our approach resulted in 10.3% and 15.3% FPS average inference rate improvements on Anomaly-ShapeNet and Real3D-AD, respectively, across a representative variety of methods. Implementation details are shown in Appendix A.6.
@@ 5.4 Generalizability Experiments and Resource Analysis
Anomaly-ShapeNet (Ours) 30.1 ($21.8\uparrow$) 2.5 ($2.6\downarrow$) 28.4 ($25.0\uparrow$) 1.3 ($1.9\downarrow$)
@@
Real3D-AD (Ours) 15.3 ($7.7\uparrow$) 4.4 ($2.4\downarrow$) 13.1 ($9.9\uparrow$) 2.9 ($2.5\downarrow$)
@@
Resource Analysis. As shown in Table 4(b), we report the resource consumption on different datasets and compare it with the model that consumes the minimum resources. Our model balances memory consumption and efficiency while maintaining SOTA performance. In the four datasets available, our rates at training and testing were improved by an average of 15.6 FPS and 13.15 FPS, respectively, compared to the most efficient models available. They only took an average memory of 3.25 GB and 2.075 GB during training and testing, respectively. This may be attributed to the fewer parameters of only 14 M, which results significant reduction compared to the PO3AD (34 M). Our model is the preferred choice in industrial anomaly detection with lower memory and a more efficient method.
5.5 Benefit of Combining External&Internal Corrective Forces.
MC4AD achieves 90.9% O-AUROC by integrating bidirectional corrective forces. Figure 4 demonstrates that external-internal forces enhance anomaly discrimination by 7.0% over single-force via complementary feature extraction mechanisms. More comparisons are shown in Appendix F.1. Our model accurately segments anomalies in complex structures while assigning lower scores to normal parts, as shown in Figure 3. More results are reported in Appendix F.2.

### v2

@@ 4.2 Corrective Force Prediction Network
Based on the Mechanics Complementary Model, we propose the Corrective Force Prediction Network (CFP-Net), which estimates internal and external corrective forces. Since defective forces constitute continuous vector fields in physical space, our differentiable U-Net architecture explicitly preserves this continuity through its topology. By U-Net’s universal approximation capability for inverse physical processes, this design directly enables learning implicit defective forces from point clouds and predicting correction forces through rigorous inverse modeling, where its differentiable operators inherently maintain continuous force field embeddings. More analysis of CTP-Net to generate damaging forces and corrective forces is reported in Appendix B.3.
@@
Three convolutional layers project $C_{1}$-dimensional features into isomorphic spaces with preserved dimensionality but distinct representations, aligning with the target feature domain.
@@
$F_{C}=MLPs(F_{9}^{\prime}).\ F_{C}\in\mathbb{R}^{n\times 6},$ (9)
@@
Anomaly Score during Inference. The norm of point-level corrective resultant force $F_{C}$ is directly considered as an anomaly score during the inference phase. This is attributed to Def. A2 in Section 3 that the model gives each point a corrective force that is a zero vector when the fact is normal and a non-zero vector when it is abnormal, as described in Eq. 4. The maximum value of the point-level score is the sample-level score.
4.3 Combined Loss
To enhance the reconstruction accuracy of corrective force fields, we propose a combined loss function that integrates three components: a symmetry loss derived from the physical constraint in Eq. 4 requiring opposing directions between internal and external correction forces in anomalous regions, a distance loss and a directional consistency loss to optimize global restoration performance jointly. The combinatorial loss function $\mathcal{L}_{comb}$ is defined:
$\mathcal{L}_{comb}$ $=\mathcal{L}_{dist}+\mathcal{L}_{dir}+\mathcal{L}_{sym}=\frac{1}{N}\sum_{i=1}^{N}\bigg{(}\frac{\|F^{\prime}_{I,i}-F^{\prime}_{E,i}\|_{1}}{\|F^{\prime}_{I,i}\|_{2}+\|F^{\prime}_{E,i}\|_{2}+\epsilon}-\frac{F^{\prime}_{I,i}}{\|F^{\prime}_{I,i}\|_{2}+\epsilon}\cdot\frac{F^{\prime}_{E,i}}{\|F^{\prime}_{E,i}\|_{2}+\epsilon}\bigg{)}$ (10)
$+\frac{1}{N}\sum_{i=1}^{N}\left\|F_{D,i}-F_{C,i}\right\|-\frac{1}{N}\sum_{i=1}^{N}\frac{F_{D,i}}{\left\|F_{D,i}\right\|_{2}+\epsilon}\cdot\frac{F_{C,i}}{\left\|F_{C,i}\right\|_{2}+\epsilon},$
@@ 4.3 Combined Loss
Quality Control (QC) under cost constraints is vital in industrial settings qc , with enterprises facing two challenges: C1: multi-level QC using coarse-to-fine detection, and C2, adaptive production of varying subcategories. We propose a Hierarchical QC (HQC) framework via three-way decision theory, combining a pruned model with our core method for C1, and introduce the Anomaly-IntraVariance dataset to address C2’s dynamic conditions. Dataset details are reported in Appendix D.
@@
$\begin{cases}\mathcal{D}_{\text{normal}}=\{P_{i}\mid\text{rank}(s_{i})\leq bN,\ s_{i}=V_{P}(P_{i})\},\\ \mathcal{D}_{\text{test}}=V_{O}(\{P_{i}\}\setminus\mathcal{D}_{\text{normal}}),\end{cases}$ (11)
@@ 5.1 Implementation
Datasets. (1) The Anomaly-ShapeNet Li_2024_CVPR has over 1,600 positive and negative samples from 40 categories, leading to a more challenging setting due to the large number of categories. (2) The Anomaly-ShapeNet-New Li_2024_CVPR provides over 450 positive and negative samples crossing 12 categories, which is not used as an evaluation dataset before. (3) The MvTec3D-AD Bergmann_2022 includes 4,147 RGB-D sample pairs from 10 categories, of which 894 are anomalous. We use the depth sample for detection evaluation. (4) The Real3D-AD liu2023real3d dataset consists of 1,254 large-scale, high-resolution point cloud samples from 12 categories, with the training set for each category only containing 4 normal samples. (5) We introduce the Anomaly-IntraVariance dataset, the first 3D dataset for simulating real-world production situations where an assembly line produces multiple products, and contains more than 1,000 samples from 16 categories. We report more details in the Appendix D.
Baselines. We selected BTF BTF , PatchCore 9879738 , M3DM m3dm , CPMF cao2023CPMF , Reg3D-AD liu2023real3d , IMRNet Li_2024_CVPR , R3D-AD r3dad , ISMP ismp and PO3AD PO3AD for comparison. We report experimental Details, how to choose baselines, and implemented baselines in Appendix A.1, A.2, and A.3, respectively.
@@
5.2 Main Experiments
Comparison on Anomaly-ShapeNet. Tables 11 and 12 provide the detection and segmentation results of different methods on Anomaly-ShapeNet across 40 categories, respectively. The proposed approach obtained an outstanding average O-AUROC and P-AUROC of 90.9% and 91.0%, outperforming the second-best methods by 7.0% and 1.2%, respectively. Moreover, our method reaches 94.0% of the average results and outperforms the second-best methods by 5.9%. Our method achieved the best average ranking on O-AUROC, P-AUROC, and O-AURP, demonstrating that our approach outperforms our competitors in most categories.
Table 1: The O-AUROC performance ($\uparrow$) of different methods on Anomaly-ShapeNet, where best and second-place results are highlighted in red and blue, respectively. The full version is in Appendix G
@@ 5.2 Main Experiments
Comparison on Real3D-AD. Tables 2 and 18 present the detection and segmentation results of different methods on Real3D-AD across 12 categories, respectively. Our model obtained O-AUROC and P-AUROC of 78.6% and 83.7%, respectively, maintaining excellent segmentation performance while achieving the SOTA detection performance. This may be because the model predicts better when the training and testing sets vary significantly.
Table 2: The O-AUROC performance ($\uparrow$) of different methods on Real3D-AD across 12 categories, where best and second-place results are highlighted in red and blue, respectively.
Comparison on Anomaly-ShapeNet-New. Tables 15 and 14 provide the detection and segmentation results of different methods on Anomaly-ShapeNet-New across 12 categories. The MC4AD obtained an outstanding average O-AUROC and P-AUROC of 88.8% and 93.7%, respectively.
@@
Comparison on Anomaly-IntraVariance. Table 19 presents the detection and segmentation results of different methods on Anomaly-IntraVariance across 16 categories, where Group 1 includes two subspecies and Group 2 has four subspecies. Our model achieves 76.1% and 80.7% on Group 1 and obtains 62.5% and 67.8% on the more challenging Group 2. The average abnormal detection metrics performed 4.40% and 2.65% better than the second-best method.
5.3 Ablation and Parameter Sensitivity Experiments
@@
O-AUROC($\uparrow$) 0.652 0.678 0.702 0.764 0.754 0.786
P-AUROC($\uparrow$) 0.688 0.705 0.733 0.812 0.792 0.837
Method O-AUROC($\uparrow$) P-AUROC($\uparrow$) FPS($\uparrow$)
$w/o_{MC-Skip}$ 0.733 0.740 17.1
@@
Analysis of Multiple Loss Functions Effectiveness.$V_{1}$” is to predict a point cloud that lacks symmetry constraints, resulting in a predicted imbalance of internal and external Corrective forces. Without $\mathcal{L}_{sym}$, the O-AUROC and P-AUROC decreased by 13.4% and 14.9%, respectively, which may be attributed to the force imbalance resulting in an incorrect offset to the normal point. Moreover, we designed “$V_{2}$”, where the model is supervised without $\mathcal{L}_{dir}$. The performance causes a degradation of the 10.8% and 13.2% anomaly detection metrics because the direction of the incorrect force is not constrained, resulting in the direction of the offset not being taken into account in the prediction of the offset. Analogous to “$V_{2}$”, “$V_{3}$” is proposed for constraining the correct size of the offset. The model performs 8.4% and 10.4% decrease in O-AUROC and P-AUROC, respectively, which may be attributed that the model meets challenges to learn the offset distance for both normal and pseudo-abnormal points without $\mathcal{L}_{dist}$. Therefore, all the proposed loss functions are crucial for predicting the defective force and thereby correctly predicting the corrective forces.
@@
Analysis of Modules Effectiveness. In “$V_{4}$”, it is observed that O-AUROC and P-AUROC are reduced by 2.2% and 2.5%, respectively, due to the replacement of our proposed anomaly generation with Norm-AS in PO3AD, which decreases the learned diversity of anomalies, and thus leads to a decrease in anomaly detection performance. In addition, our pruned version of “$V_{P}$” acquires a more efficient 20 FPS (>14.8 FPS) while maintaining excellent anomaly detection performance, making our HQC strategy potentially possible, as discussed in Section 5.4. Moreover, the model without MC-Skip ($w/o\ MC-Skip$) modules failed to capture the complementary features of external and internal corrective forces and was not able to generate the resultant corrective force, resulting in a decrease in the anomaly detection metrics by 5.3% and 9.7%, respectively. In contrast, the model directly outputs 3-channel offsets ($Out=3$) equivalent to the direct learning of the corrective force, resulting in a decrease of 4.0% and 4.2% in P-AUROC and O-AUROC, respectively. Layer evaluations revealed that reducing layers from 8 to 2 decreased anomaly detection metrics by 5.8% and 12.4%, prompting the selection of 8 layers for performance-efficiency balance. More parameter sensitivity experiments are reported in the Appendix E.
5.4 Generalizability Experiments and Resource Analysis
Generalizability Experiments. We apply MC-Skip to the time-tested classical model, including PointNet pointnet , PointNet++ pointnet2 , PointTransformer pointtransformer , and DGCNN dgcnn , to validate the generalizability of our model. As shown in Table 5(a), the accuracies with and without normal participation on the classification task wu20153dshapenetsdeeprepresentation increased on average by 0.65% and 0.6%, respectively, and the instance average IoU and class average IoU on the segmentation task shapenet increased on average by 0.775% and 0.675%, respectively. Moreover, our HQC strategy performs well on existing 3D anomaly detection models and generalizes to memory bank, as reported in Table 4. Our approach resulted in 10.3% and 15.3% FPS average inference rate improvements on Anomaly-ShapeNet and Real3D-AD, respectively, across a representative variety of methods. Implementation details are shown in Appendix A.6.
Table 4: Comparison between one model and our HQC.
@@ 5.4 Generalizability Experiments and Resource Analysis
One Model HQC One Model HQC
O-AUROC($\uparrow$) P-AUROC($\uparrow$) FPS($\uparrow$) O-AUROC($\uparrow$) P-AUROC($\uparrow$) FPS($\uparrow$) O-AUROC($\uparrow$) P-AUROC($\uparrow$) FPS($\uparrow$) O-AUROC($\uparrow$) P-AUROC($\uparrow$) FPS($\uparrow$)
@@
Anomaly-ShapeNet (Ours) 30.1 ($21.8\uparrow$) 2.5 ($2.6\downarrow$) 25.0 ($21.6\uparrow$) 1.3 ($1.9\downarrow$)
@@
Real3D-AD (Ours) 15.3 ($7.7\uparrow$) 4.4 ($2.4\downarrow$) 14.8 ($11.6\uparrow$) 2.9 ($2.5\downarrow$)
@@
Resource Analysis. As shown in Table 5(b), we report the resource consumption on different datasets and compare it with the model that consumes the minimum resources. In the four datasets available, our rates at training and testing were improved by an average of 15.6 FPS and 12.725 FPS, respectively, compared to the most efficient models available. They only took an average memory of 2.75 GB and 1.075 GB during training and testing, respectively. This may be attributed to the fewer parameters of only 14 M, which results in a significant reduction compared to the PO3AD (34 M). Our model is the preferred choice in industrial anomaly detection, as it is a more efficient method with lower memory.
@@
Figure 4: Localization results.
Benefit of Combining External & Internal Corrective Forces. MC4AD achieves 90.9% O-AUROC by integrating bidirectional corrective forces. Figure 3 demonstrates that external-internal forces enhance anomaly discrimination by 7.0% over single-force via complementary feature extraction mechanisms. More comparisons are shown in Appendix F.1. Our model accurately segments anomalies in complex structures while assigning lower scores to normal parts, as shown in Figure 4. More results are reported in Appendix F.2.